Trigonometry

the number k is such that tan{arc tan(2)+arc tan(20k)}=k.

the sum of allpossible values of k is:

sonal nagwani
14 Years agoGrade

1 Answer

vignesh prasad
14 Years ago

 tan{arc tan(2)+arc tan(20k)}=k.

using formula,arctan x + arctan y = arctan ((x+y)/(1-xy))

we get,

(20k+2)/(1-40k)=k

on solving we get ,

40k^2+19k+2=0,

=>sum=-19/40